OpenStudy (anonymous):

What is the slope of the equation 5x + 4y = 8 ?

OpenStudy (anonymous):

5x + 4y = 8 -5x -5x 4y = -5x + 8 /4 /4 y = -5/4x +2 slope = -5/4

OpenStudy (anonymous):

The slope = (2-0)/(0-8/5) = - 10/8 = - 5/4